# Challenges and Opportunities for Visual Analytics in Jurisprudence

### v2

@@ 3 Understanding Legal Research Practice
Jurisprudence is a cornerstone of the legal domain, yet its complexity and reliance on vast information repositories pose significant challenges. A deep understanding of legal research practices is essential to leverage VA to build effective applications. To this end, we need to establish a common understanding of how legal scholars operate and which tasks are relevant to them. While there is extensive and foundational work on how jurisprudence operates [77, 78, 79, 49], our work contributes new insights from the perspective of computer science by synthesizing existing knowledge and incorporating findings from the interviews with legal scholars.
3.1 How Legal Scholars Operate
@@
The German civil law system builds upon a codified framework, anchored in the Grundgesetz für die Bundesrepublik Deutschland (Basic Law) and several codes, such as the Bürgerliches Gesetzbuch (Civil Code). These codes form the foundation of the legal system and are supplemented by an extensive array of additional legal sources. Until the end of 2024, there are over 90,000 legal norms and regulations at the federal level, supplemented by numerous state and local statutes [91]. Beyond codified laws, the system includes other essential sources such as explanatory memoranda, legal commentaries, court rulings, and scholarly publications in trade journals and commemorative volumes [49, p. 52]. Explanatory memoranda, published by the German parliament, elaborate on the legislative intent and provide interpretive guidance [49, p. 55]. Legal commentaries, authored by scholars, interpret and analyze legal norms, while court rulings refine their application in practice [49, p. 52]. However, access to these sources is often restricted. In Germany, for example, legal commentaries are primarily available on the proprietary platforms beck-online [26] and juris [27], which impose licensing constraints. Similarly, court rulings – particularly from lower courts – are frequently inaccessible due to data protection regulations. While some publicly available datasets exist (refer to Table 1), they represent less than 1% of legal decisions [25, 18, 24].
@@
Systematic Relationships — The interplay of these diverse legal sources constitutes the fabric of the German legal system [4, p. 460]. Understanding individual components requires knowledge of their relationships within the broader system. These relationships are formally structured, hierarchical, and semantic, necessitating tacit knowledge to navigate their complexity effectively. Applying legal sources follows a systematic process that relies on explicit procedures [4, p. 412]. Before applying legal norms, practitioners must first understand the sources that constitute the law. Then, the application process typically unfolds through a series of distinct steps [4, p. 414]. First, practitioners establish the facts of the case, carefully analyzing the situation to identify relevant details. Experienced practitioners often intuitively classify the case within a specific legal area or discipline [4, p. 413]. A particularly challenging aspect of this process involves identifying the legal norms relevant to the case. This task requires understanding the formally structured, hierarchical, and semantic relationships within the legal system [4, p. 460]. Interpreting these legal texts depends on contextualizing a given norm within its broader legal framework, as its meaning and implications are frequently defined by its connections to others. Determining these connections relies heavily on the practitioner’s expertise and familiarity with the legal system. Practitioners must continuously evaluate these relationships, considering both explicit legal provisions and implicit analogies or concepts derived from prior decisions and academic literature [4, p. 413]. After identifying the relevant norms, practitioners assess whether the case fulfills the specific elements outlined by these norms. This process, known as subsumption, involves comparing the facts of the case with the requirements of the legal provisions [4, p. 413]. If the elements are satisfied, the practitioner determines the legal consequences accordingly. However, cases are often not straightforward, and legal scholars can have differing opinions.
Nevertheless, through this systematic and knowledge-intensive process, practitioners bridge the gap between abstract legal principles and real-world cases, underscoring the critical role of domain expertise in legal reasoning. While many professionals engage with this legal system, our work focuses on jurisprudence, the scientific debate and analysis of law. However, the tasks and workflow described in the following also apply broadly to other legal professions.
3.2 Tasks
Legal scholars engage in tasks that span multiple levels of complexity, which can be categorized into fundamental information retrieval, understanding structural relationships, advancing legal reasoning, and specialized tasks.
@@ 3.2 Tasks
Fundamental Information Retrieval ($T_{1}$) — Scholars focus on identifying and retrieving relevant legal sources at the foundational level. These tasks include filtering sources pertinent to a particular topic [76], understanding the formal legal organization of these sources, and determining the location of specific legal norms within the broader legal corpus [75]. Efficient information retrieval serves as the groundwork for deeper analytical tasks.
@@
Understanding Structural Relationships ($T_{2}$) — Building on information retrieval, scholars analyze the structural relationships within legal texts. The analysis involves evaluating the chronological order of statutes [76] and interpreting both explicit and implicit relationships between legal norms [74, 76]. Grasping these relationships is critical for understanding how individual legal norms interact within the broader legal framework, enabling scholars to navigate the complexities of formal legal structures effectively.
@@
Advancing Legal Reasoning ($T_{3}$) — At the highest level, legal reasoning tasks require synthesizing information to uncover patterns and identify conflicts across legal texts [70, 67]. Scholars engage in activities such as detecting inconsistencies between norms, identifying legal precedents, and formulating arguments. Additionally, they analyze legal conflicts [92] and derive insights that inform decision-making processes.
@@
Specialized Tasks ($T_{S}$) — Beyond these general categories, certain tasks are unique to specific legal disciplines or application scenarios. Examples include Policy Modeling as analyzing the implications of proposed laws and policies [92], resolving natural language queries through interpreting and answering complex legal queries [75] or comparative law analysis which involves comparing legal systems or statutes across jurisdictions [74].
@@ 4.1 Methodology
The semi-structured interviews with subject matter experts ($n=9$) from the legal domain, averaging one hour each, followed a protocol but allowed for follow-ups to capture detailed insights. We conducted the interviews at the participants’ workplace. Each interview began by gathering information about the participants’ specializations and backgrounds, asking them about their legal expertise as well as their professional and academic experiences. Next, we inquired about the established workflows for working with legal texts in each of these specializations, including their tasks. The participants used their familiar setup to demonstrate how they practice their workflows. We asked them to showcase these workflows using recent examples of their research. We also wanted the participants to describe the tools they use and their role in their workflows. We also encouraged participants to compare their practices in previous roles to establish diverse tasks. Finally, we asked the experts to identify specific obstacles they face throughout their workflows with the tools they described, providing examples. Looking forward, we inquired about their needs and expectations for applications and tools that could improve their workflows.
@@ 4.2 Participants
The selection of experts represents a diverse range of experiences: The participants included a professor of public law ($E_{1}$), seven post-doctoral ($E_{2}$) and doctoral ($E_{3-8}$) research associates specializing in various legal areas, and an undergraduate student of public management ($E_{9}$). Concerning their experience, $E_{1}$ has extensive experience in jurisprudence and legal practice, having spent 20 years in academia. One set of participants focuses on German federal ($E_{3}$), state public ($E_{4}$), and European ($E_{8}$) law, and they all have practical experience in legal institutions (e.g., public prosecutor offices, research services of the German parliament, European Commission, legal committees, the German ministry of the interior). The remaining participants specializes in German civil law ($E_{2}$), administrative law ($E_{5}$), criminal law ($E_{6}$), private building law ($E_{7}$), while $E_{9}$ focuses on local politics and municipal law. This diversity ensures a comprehensive understanding of workflows across legal contexts.
@@ 4.3 Interviews
Scholars follow a structured yet often non-linear approach for this task. For example, $E_{7}$ succinctly summarizes their approach by identifying the law, understanding its structure, reading the reasoning, and finally delving into the jurisdiction and literature. $E_{3}$, instead, initially conducts a keyword search to identify relevant legal norms. Other participants ($E_{2}$, $E_{4}$, $E_{5}$, $E_{7}$, and $E_{8}$) concur and emphasize the significance of using keyword search tools. All participants consistently refer to beck-online and juris for this task, which they consider their primary literature database. However, for European legislation, $E_{3}$ also mentions using Google Search, while $E_{4}$ refers to university document collections. The variety of legal databases highlights a problem with the current user interfaces (UI), which are often incomplete silos that do not connect different databases, in particular when studying law across different legislative systems, hindering efficient workflows ($E_{2}$). Some experts also wish for interface elements beyond the text that show interrelations and dependencies of laws that block specific processes ($E_{1}$), or how legal processes and administrative procedures follow step-by-step ($E_{1}$ and $E_{7}$). Regarding beck-online and juris, some scholars express concern, noting that ”keyword search is always quite poor, often with zero hits because [it is] too specific”  ($E_{4}$). On the other hand, for too broad queries, $E_{2}$ frequently finds themselves manually screening through search results using their browser’s search functionality. $E_{6}$ also expresses dissatisfaction with the inefficiency of these keyword searches. $E_{3}$, $E_{4}$, $E_{8}$, and $E_{9}$ highlight the manual aspect of their workflow when skimming through the table of contents of laws to locate potentially relevant legal norms.
@@
At the same time, the participants acknowledge that ”many things are not in the text”  ($E_{3}$) but require ”the activation of implicit knowledge”  ($E_{6}$). However, they regret that ”search engines do not have this context”  ($E_{6}$). $E_{4}$ further emphasizes the interconnected nature of legal materials, highlighting their frequent reliance on explanatory memoranda, commentary, and court rulings that provide context for legal norms. As $E_{3}$ confirms, ”detailed commentary is more interesting than shorter discussions.” $E_{5}$ also underscores the significance of legal materials linked to other laws and legal norms through references, which facilitate systematic interpretation. As $E_{6}$ puts it, ”legal thinking is highly structured.” In light of this, $E_{3}$ and $E_{7}$ advocate for abstracting legal norms into concepts that enhance their research and facilitate compatibility with other legal systems. For instance, $E_{4}$ employs concepts to compare court rulings that are not thematically related but contain legal ”reasoning and evidence” relevant to their discipline. Furthermore, the meaning of terms can be defined elsewhere, leading to ”definition cascades”  ($E_{3}$) that require resolution through multiple layers and across different legal texts. A related problem is the collision of disparate legal fields and competing legal definitions in intersecting areas ($E_{1}$, $E_{9}$), in particular, if criteria like lex specialis (i.e., a specific law takes precedence over a general law) or lex posterior (i.e., a newer law takes precedence over older legislation) cannot be applied, or a balancing of interests is required. Despite these considerations, legal scholars must remain vigilant in wording, as $E_{1}$ emphasizes, ”three words change and the case looks completely different.” Further, the law only provides the guardrails and general pathways, as the legal interpretation allows for some discretionary decisions depending on the specific cases ($E_{6}$), for example, in agencies ($E_{9}$). Legal scholars must bridge the gap between their domain knowledge and perspective, also communicating with their stakeholders, including employees, managers, politicians, and the general public, in their interpretation ($E_{1}$).
@@
In the face of these requirements, participants remain skeptical about digitization and LLMs in the legal domain. First and foremost, $E_{8}$ continues to identify ”digitization as a problem” within the legal domain. For instance, electronic files in courts still encounter practical challenges ($E_{8}$), and court rulings are often not digitally accessible ($E_{4}$) or follow restrictive copyright and licensing models ($E_{6})$. $E_{4}$ further highlights that these inaccessible documents disrupt their workflow, as they must scan them before proceeding with their research. Regarding LLMs, $E_{5}$ finds that using GPT still requires a lot of trial and error to achieve satisfactory results when connected with search engines. $E_{2}$ agrees that LLMs ”lack traceability” but still believes that AI could be relevant in supporting whether a text is relevant and worthwhile to consider during legal research. However, participants still prefer to read them in full themselves, which $E_{8}$ also finds to facilitate serendipitous discovery of relevant legal norms and arguments. $E_{7}$ occasionally uses LLMs for mundane tasks only tangibly to legal research questions, such as preparing interviews.
@@ 5 Addressing Challenges in Legal Research: A Three-Phase Workflow Design
Despite the digitization of law and recent advancements in AI, our interviews (refer to subsection 4.3) reveal challenges that cannot be resolved solely through technological solutions. The research process of legal scholars is fraught with difficulties, ranging from accessing relevant legal materials amid restrictive license models to navigating formal legal structures and relationships between legal documents. Moreover, synthesizing tacit knowledge into transparent and well-reasoned arguments poses a significant challenge. A syntactic keyword search in commercial databases often proves inadequate to capture these nuanced interpretation tasks, highlighting the limitations of current UIs and underscoring the need for support contextualized to the user.
In line with these insights, and building on the bottom-up findings from our subject-matter expert interviews, we observe that legal scholars follow structured steps in their research [4, pp. 412–413]. These phases require different types of support. While scholars typically switch between these phases effortlessly, addressing the need for support and the difficulties of the legal research process from a technical perspective necessitates a coherent structure to the workflow of legal scholars. This allows for using different support mechanisms in each phase [93]. Most participants begin by conducting a keyword search to identify relevant laws and legal norms. However, a single attempt is usually insufficient for satisfactory results, so keyword search becomes an iterative process [4, p. 413]. It serves as a tool to facilitate the analysis of legal materials, but it is not an end in itself. Often, key insights emerge not through targeted searches but during comprehensive reading. Subsequently, legal scholars transition to contextualizing laws and legal norms with each other, using commentary and other legal materials [4, pp. 93, 459–461]. This involves connecting the initially narrow scope of documents with a broader landscape of legal materials to enable systematic interpretation. Often, identifying individual legal norms is not sufficient, but connecting them to legal concepts is necessary to facilitate legal reasoning. Given the vast legal document collections, this process can be cognitively taxing for scholars. Therefore, they rely on documentation throughout the process [66]. Documenting their legal insights enables synthesis, which is used to construct legal arguments and constitute legal knowledge. While many of the recurring challenges faced by researchers arise during their work, some are more prevalent at specific phases. Based on these observations and the literature, we identify a three-phase workflow that consists of an initial overview, a strategic analysis and contextualization, and synthesis.
5.1 Three-Phase Workflow
@@
Figure 1: The integration of jurisprudence and Visual Analytics (VA) for personalized legal research: This approach processes diverse legal documents – including laws, explanatory memoranda, and court rulings – to support legal scholars in three key tasks: (1) fundamental information retrieval, (2) understanding structural relationships, and (3) advancing legal reasoning, while addressing several challenges. Through interviews with domain experts, we identify three workflow phases, namely (1) Discovery & Scoping, (2) Analysis & Interpretation, and (3) Synthesis & Documentation. Through them, scholars iteratively navigate through and generate analytical artifacts that capture explicit and tacit knowledge, fostering deeper comprehension of legal structures and reasoning, and informing a VA design for jurisprudence.
This three-phase workflow ($W_{\texttt{P1}}$- $W_{\texttt{P3}}$) is frequently encountered by legal scholars (see also Figure 1). The phases that we define – Discovery & Scoping ($T_{1}$), Analysis & Interpretation ($T_{2}$), and Synthesis & Documentation ($T_{3}$) – illustrate the systematic but manual processes scholars typically follow. To illustrate, consider Alice, a legal scholar with extensive experience in German legislation. She is researching racial profiling in the context of the legal safeguards offered by German law against law enforcement from engaging in racial profiling. The following three sections describe the workflow’s phases as executed traditionally without VA, outlining legal scholars’ challenges, where $E_{i}$ refers to an exemplary experience described by a participant in section 4.
5.1.1 $W_{\texttt{P1}}$: Discovery and Scoping
[…]

### v3

@@ 3 Understanding Legal Research Practice
Jurisprudence is a cornerstone of the legal domain, yet its complexity and reliance on vast information repositories pose significant challenges. Understanding legal research practices is essential to building effective applications with VA. Hence, we need to establish a common understanding of how legal scholars operate and which tasks are relevant to them. While jurisprudence is described extensively [borchardJurisprudenceGermany1912, brugger1994legal, foster2010germanlaw, bundesministeriumfuerjustizHandbuchRechtsfoermlichkeit2024], our work contributes new insights from the perspective of VA, incorporating existing knowledge and findings from the interviews with legal scholars.
3.1 How Legal Scholars Operate - A Technical Perspective
@@
The German civil law system builds upon a codified framework, anchored in the Grundgesetz für die Bundesrepublik Deutschland (Basic Law) and several codes, such as the Bürgerliches Gesetzbuch (Civil Code). These codes form the foundation of the legal system and are supplemented by an extensive array of additional legal sources. Until the end of 2024, there are over 90,000 legal norms and regulations at the federal level, supplemented by numerous state and local statutes [mwo1797GesetzeUnd]. Beyond codified laws, the system includes other essential sources such as explanatory memoranda, legal commentaries, court rulings, and scholarly publications in trade journals and commemorative volumes [bundesministeriumfuerjustizHandbuchRechtsfoermlichkeit2024, p. 52]. Explanatory memoranda, published by the German parliament, elaborate on the legislative intent and provide interpretive guidance [bundesministeriumfuerjustizHandbuchRechtsfoermlichkeit2024, p. 55]. Legal commentaries, authored by scholars, interpret and analyze legal norms, while court rulings refine their application in practice [bundesministeriumfuerjustizHandbuchRechtsfoermlichkeit2024, p. 52]. However, access to these sources is often restricted. In Germany, for example, legal commentaries are primarily available on the proprietary platforms beck-online [verlagc.h.beckohgBeckonline] and juris [jurisgmbhJuris], which impose licensing constraints. Similarly, court rulings – particularly from lower courts – are frequently inaccessible due to data protection regulations. While some publicly available datasets exist (refer to 1), they represent less than 1% of legal decisions [ltoTransparenzJustizStagnation, vogelRichterKeinBot2024, zanderUrteileUnterVerschluss2024].
@@
Systematic Relationships — The interplay of these diverse legal sources constitutes the fabric of the German legal system [ruethersRechtstheorieUndJuristische2022, p. 460]. Understanding individual components requires knowledge of their relationships within the broader system. These relationships are formally structured, hierarchical, and semantic, necessitating tacit knowledge to navigate their complexity effectively. Applying legal sources follows a systematic process that relies on explicit procedures [ruethersRechtstheorieUndJuristische2022, p. 412]. Before applying legal norms, practitioners must first understand the sources that constitute the law. Then, the application process typically unfolds through a series of distinct steps [ruethersRechtstheorieUndJuristische2022, p. 414]. First, practitioners establish the facts of the case, carefully analyzing the situation to identify relevant details. Experienced practitioners often intuitively classify the case within a specific legal area or discipline [ruethersRechtstheorieUndJuristische2022, p. 413]. A challenging aspect of this process involves identifying the legal norms relevant to the case. This task requires understanding the formally structured, hierarchical, and semantic relationships within the legal system [ruethersRechtstheorieUndJuristische2022, p. 460]. Interpretation depends on contextualizing a norm within its broader legal framework, as its meaning and implications are frequently defined by its context. Determining these connections relies heavily on the practitioner’s expertise and familiarity. Practitioners must continuously evaluate these relationships, considering both explicit legal provisions and implicit concepts derived from prior decisions and academic literature [ruethersRechtstheorieUndJuristische2022, p. 413]. After identifying the relevant norms, practitioners assess whether the case fulfills the specific elements outlined by these norms. This process, known as subsumption, involves comparing the facts with the requirements of the legal provisions [ruethersRechtstheorieUndJuristische2022, p. 413]. If the elements are satisfied, the practitioner determines the legal consequences accordingly. However, cases are often not straightforward, and legal scholars can have differing opinions.
3.2 High-Level Analysis Task Categories
Legal tasks that span multiple levels of complexity, which can be categorized into the following, which we carefully base on the literature [bundesministeriumfuerjustizHandbuchRechtsfoermlichkeit2024, ruethersRechtstheorieUndJuristische2022, lacavaLawNetVizWebbasedSystem2022, lettieriLegalMacroscopeExperimenting2017, resckLegalVisExploringInferring2023, gomez2015understanding, mclachlanVisualisationLawLegal2021, burkhardtVisualLegalAnalytics2019] and our previous findings in 3.1. For the process of distilling task categories, we orient ourself at the principles established in the HCI community [sedlmairDesignStudyMethodology2012, lamEmpiricalStudiesInformation2012, brehmerMultiLevelTypologyAbstract2013] to identifying, categorizing, grouping, and abstracting tasks:
@@ 3.2 High-Level Analysis Task Categories
Fundamental Information Retrieval ($T_{1}$) — Scholars focus on identifying and retrieving relevant legal sources at the foundational level, which includes filtering sources by topic [resckLegalVisExploringInferring2023], understanding the formal legal organization, and determining specific legal norms within the broader legal corpus [lacavaLawNetVizWebbasedSystem2022]. This includes downstream tasks such as finding the location of a legal norm [lacavaLawNetVizWebbasedSystem2022].
@@
Understanding Structural Relationships ($T_{2}$) — Building on information retrieval, scholars analyze the structural relationships within legal texts, which involves evaluating the order of statutes [resckLegalVisExploringInferring2023] and interpreting both explicit and implicit relationships [lettieriLegalMacroscopeExperimenting2017, resckLegalVisExploringInferring2023]. Grasping these relationships is critical for understanding how individual legal norms interact within the broader legal framework. This category includes downstream tasks such as recognizing a lex specialis for a general law.
@@
Advancing Legal Reasoning ($T_{3}$) — At the highest level, legal reasoning tasks require synthesizing information to uncover patterns and identify conflicts across legal texts [gomez2015understanding, mclachlanVisualisationLawLegal2021]. Scholars engage in activities such as detecting inconsistencies between norms, identifying legal precedents, and formulating arguments. Additionally, they analyze legal conflicts [burkhardtVisualLegalAnalytics2019] and derive insights informing decision-making processes. This includes downstream tasks such as reasoning about flaws and regulatory gaps [ruethersRechtstheorieUndJuristische2022].
@@
Specialized Tasks ($T_{S}$) — Beyond these general categories, certain tasks are unique to specific legal disciplines or application scenarios. Examples include Policy Modeling such as analyzing the implications of proposed laws and policies [burkhardtVisualLegalAnalytics2019], resolving natural language queries through interpreting and answering complex legal queries [lacavaLawNetVizWebbasedSystem2022] or comparative law analysis, which compares jurisdictions [lettieriLegalMacroscopeExperimenting2017].
@@ 4.1 Methodology
The semi-structured interviews with subject matter experts ($n=9$) from the legal domain, averaging one hour each, followed a protocol but allowed for follow-ups to capture detailed insights. We conducted the interviews at the participants’ workplace. Each interview began by gathering information about the participants’ specializations and backgrounds, asking them about their legal expertise, and professional and academic experiences. Next, we inquired about the established workflows for working with legal texts in specialization, including their tasks. The participants used their familiar setup to demonstrate how they carry out their workflows. We asked them to showcase these workflows using recent examples of their research. We also wanted the participants to describe the tools they use and their role in their workflows. We also encouraged participants to compare their practices in previous roles to establish diverse tasks. Finally, we asked the experts to identify specific obstacles they face with the tools they described, providing examples. Looking forward, we inquired about their needs and expectations for tools that could improve their workflows.
@@ 4.2 Participants
The selected experts represents a diverse range of experiences: The participants included a professor of public law ($E_{1}$), seven post-doctoral ($E_{2}$) and doctoral ($E_{3-8}$) research associates specializing in various legal areas, and an undergraduate student of public management ($E_{9}$). Concerning their experience, $E_{1}$ has experience in jurisprudence and legal practice, having spent 20 years in academia. One set of participants focuses on German federal ($E_{3}$), state public ($E_{4}$), and European ($E_{8}$) law, all having practical experience in legal institutions (e.g., public prosecutor offices, research services of the German parliament, European Commission, legal committees, the German ministry of the interior). The remaining participants specialize in German civil law ($E_{2}$), administrative law ($E_{5}$), criminal law ($E_{6}$), private building law ($E_{7}$), while $E_{9}$ focuses on local politics and municipal law.
@@ 4.3 Interviews
Scholars follow a structured yet often non-linear approach for this task. For example, $E_{7}$ succinctly summarizes their approach by identifying the law, understanding its structure, reading the reasoning, and finally delving into the jurisprudence and literature. $E_{3}$, instead, initially conducts a keyword search to identify relevant legal norms. Other participants ($E_{2}$, $E_{4}$, $E_{5}$, $E_{7}$, and $E_{8}$) concur and emphasize the significance of using keyword search tools. All participants consistently refer to beck-online and juris for this task, which they consider their primary literature database. However, for European legislation, $E_{3}$ also mentions using Google Search, while $E_{4}$ refers to university document collections. The variety of legal databases highlights a problem with the current user interfaces (UI), which are often incomplete silos that do not connect different databases, in particular when studying law across different legislative systems, hindering efficient workflows ($E_{2}$). Some experts also wish for interface elements beyond the text that show interrelations and dependencies of laws that block specific processes ($E_{1}$), or how legal processes and administrative procedures follow step-by-step ($E_{1}$ and $E_{7}$). Regarding beck-online and juris, some scholars express concern, noting that “keyword search is always quite poor, often with zero hits because [it is] too specific” ($E_{4}$). On the other hand, for too broad queries, $E_{2}$ frequently finds themselves manually screening through search results using their browser’s search functionality. $E_{6}$ also expresses dissatisfaction with the inefficiency of these keyword searches. $E_{3}$, $E_{4}$, $E_{8}$, and $E_{9}$ highlight the manual aspect of their workflow when skimming through the table of contents of laws to locate potentially relevant legal norms.
@@
At the same time, the participants acknowledge that “many things are not in the text”  ($E_{3}$) but require “the activation of implicit knowledge” ($E_{6}$). However, they regret that “search engines do not have this context” ($E_{6}$). $E_{4}$ further emphasizes the interconnected nature of legal materials, highlighting their frequent reliance on explanatory memoranda, commentary, and court rulings that provide context for legal norms. As $E_{3}$ confirms, “detailed commentary is more interesting than shorter discussions.” $E_{5}$ also underscores the significance of legal materials linked to other laws and legal norms through references, which facilitate systematic interpretation. As $E_{6}$ puts it, “legal thinking is highly structured.” In light of this, $E_{3}$ and $E_{7}$ advocate for abstracting legal norms into concepts that enhance their research and facilitate compatibility with other legal systems. For instance, $E_{4}$ employs concepts to compare court rulings that are not thematically related but contain legal “reasoning and evidence” relevant to their discipline. Furthermore, the meaning of terms can be defined elsewhere, leading to “definition cascades” ($E_{3}$) that require resolution through multiple layers and across different legal texts. A related problem is the collision of disparate legal fields and competing legal definitions in intersecting areas ($E_{1}$, $E_{9}$), in particular, if criteria like lex specialis (i.e., a specific law takes precedence over a general law) or lex posterior (i.e., a newer law takes precedence over older legislation) cannot be applied, or a balancing of interests is required. Despite these considerations, legal scholars must remain vigilant in wording, as $E_{1}$ emphasizes, “three words change and the case looks completely different.” Further, the law only provides the guardrails and general pathways, as the legal interpretation allows for some discretionary decisions depending on the specific cases ($E_{6}$), for example, in agencies ($E_{9}$). Legal scholars must bridge the gap between their domain knowledge and perspective, also communicating with their stakeholders, including employees, managers, politicians, and the general public, in their interpretation ($E_{1}$).
@@
In the face of these requirements, participants remain skeptical about digitization and LLMs in the legal domain. First and foremost, $E_{8}$ continues to identify “digitization as a problem” within the legal domain. For instance, electronic files in courts still encounter practical challenges ($E_{8}$), and court rulings are often not digitally accessible ($E_{4}$) or follow restrictive copyright and licensing models ($E_{6})$. $E_{4}$ further highlights that these inaccessible documents disrupt their workflow, as they must scan them before proceeding with their research. Regarding LLMs, $E_{5}$ finds that using GPT still requires a lot of trial and error to achieve satisfactory results when connected with search engines. $E_{2}$ agrees that LLMs “lack traceability” but still believes that AI could be relevant in supporting whether a text is relevant and worthwhile to consider during legal research. However, participants still prefer to read them in full themselves, which $E_{8}$ also finds to facilitate serendipitous discovery of relevant legal norms and arguments. $E_{7}$ occasionally uses LLMs for mundane tasks only tangibly to legal research questions, such as preparing interviews.
@@ 5 Addressing Challenges in Legal Research: A Three-Phase Workflow Design
Figure 1: Using Visual Analytics (VA) for legal research can simplify working with diverse legal documents – including laws, explanatory memoranda, and court rulings – to support legal scholars in three key tasks: (1) fundamental information retrieval, (2) understanding structural relationships, and (3) advancing legal reasoning, while addressing several challenges. Through interviews with domain experts, we identify three workflow phases, namely (1) Discovery & Scoping, (2) Analysis & Interpretation, and (3) Synthesis & Documentation. Through them, scholars iteratively navigate through and generate analytical artifacts that capture explicit and tacit knowledge, fostering a deeper comprehension of legal structures and reasoning, and informing a VA design for jurisprudence.
In line with these insights, and building on the bottom-up findings from our subject-matter expert interviews, we observe that legal scholars follow structured steps in their research [ruethersRechtstheorieUndJuristische2022, pp. 412–413]. Most participants begin by conducting a keyword search to identify relevant laws and legal norms. However, a single attempt is usually insufficient for satisfactory results, so keyword search becomes an iterative process [ruethersRechtstheorieUndJuristische2022, p. 413]. Often, key insights emerge not through targeted searches but during comprehensive reading. Subsequently, legal scholars transition to contextualizing laws and legal norms with each other, using commentary and other legal materials [ruethersRechtstheorieUndJuristische2022, pp. 93, 459–461]. This involves connecting the initially narrow scope of documents with a broader landscape of legal materials to enable systematic interpretation. Given the vast legal document collections, this process can be cognitively taxing for scholars. Therefore, they rely on documentation throughout the process [cantatoreMakingConnectionsIncorporating2016]. Documenting their legal insights enables synthesis, which is used to construct legal arguments and constitute legal knowledge. Based on these observations and the literature, we identify a three-phase workflow ($W_{\texttt{P1}}$- $W_{\texttt{P3}}$) (see also 1). The phases that we define – Discovery & Scoping ($T_{1}$), Analysis & Interpretation ($T_{2}$), and Synthesis & Documentation ($T_{3}$) – illustrate the systematic but manual processes scholars typically follow. To illustrate, consider Alice, a legal scholar with extensive experience in German legislation. She is researching racial profiling in the context of the legal safeguards offered by German law against law enforcement from engaging in racial profiling. The following three sections describe the workflow’s phases as executed traditionally without VA, outlining legal scholars’ challenges, where $E_{i}$ refers to an exemplary experience described by a participant in 4.
5.1 $W_{\texttt{P1}}$: Discovery & Scoping
In the first phase of the workflow ($W_{\texttt{P1}}$), Alice identifies suitable legal texts and narrows them down to address her research problem ($E_{7}$$T_{1}$). Familiar with the commercial legal database beck-online [verlagc.h.beckohgBeckonline], she begins by querying the platform’s search interface. Alice knows from recurring media reports that cases of racial profiling are sometimes prevalent in operations conducted by the Bundespolizei (German Federal Police), whose responsibilities include securing German borders (§ 2 BPolG (Federal Police Act)), train stations (§ 3 BPolG), and airports (§ 4 BPolG) (Reliance on Tacit Knowledge). She starts with the search phrase “Racial Profiling Bundespolizei.” Her query yields only five matching documents, prompting Alice to broaden her search ($E_{8}$) using the more general phrase “Racial Profiling.” While this query produces a significantly larger volume of results, it also introduces many irrelevant entries (Inefficiency of Current Interfaces). Alice begins her review by examining the most relevant legal norms by beck-online for her search queries. These include Art. 3 GG (Basic Law), § 23 BPolG, and Art. 14 EMRK (European Convention on Human Rights).